# Preference Curriculum: LLMs Should Always Be Pretrained on Their Preferred Data

### v1

@@ 3.2 Main Results
Sequential PD $\downarrow$ 54.7 21.8 87.1 33.7 67.8 53.0
PD $\uparrow$ 56.1 21.3 86.2 34.4 67.6 53.1
PPL $\downarrow$ 45.5 20.6 71.2 30.3 63.7 46.3
PPL $\uparrow$ 47.8 17.9 72.7 29.1 62.4 46.0
Qu.Edu $\downarrow$ 57.2 26.4 85.4 33.0 66.2 53.6
Qu.Edu $\uparrow$ 56.8 26.0 84.1 33.5 67.9 53.7
Preference CL PPL S.R. 56.1 24.1 87.8 33.9 67.4 53.9
PPL S. 56.1 22.6 85.5 34.2 67.5 53.2
Qu.Edu S.R 56.7 24.9 86.2 33.6 66.9 53.7
Qu.Edu S. 55.5 24.8 87.8 34.0 67.4 53.9
PD S.R. 56.7 24.9 86.2 33.6 67.4 53.8
PDPC PD S. 57.3 26.6 87.9 33.7 68.0 54.7
Table 1: Downstream tasks results for different setings on 1.3B models. We report accuracy for each task, and the best performances are marked in bold. Abbreviations: HellaSw. = HellaSwag, AVG. = Average, S.=S-shape Function, S.R.=S-shape Reverse Function
@@
PDPC PD S. 69.7 35.8 95.3 76.3 35.8 35.6 36.1 54.9
Table 2: Downstream tasks results for different setings after training 3B models on 1T tokens.
@@
PDPC with $n=2$ consistently outperforms all baselines, regardless of the metric used, showing significant performance and convergence improvements over the baseline. Notably, the 3B model trained on 1T tokens with PDPC demonstrates an average accuracy increase of 4.1% across all benchmarks and 8.1% across MMLU and CMMLU, highlighting the effectiveness of our framework. Figure 3 depicts performance improvements in the 1.3B and 3B models as training progresses, with our method significantly outperforming Random in the latter half of pre-training. In this phase, data dominated by high PD is crucial, especially in the 3B model, highlighting the effectiveness of transitioning from small to large PD data, which significantly boosts model performance and promotes emergent capabilities.
@@
[…]
@@
Comparison of the results from ’Preference CL, PPL, S.’, ’Preference CL, Qu.Edu, S.’, and ’PDPC, PD, S.’ shows that using PD as a metric yields the best results, particularly on the ARC-C and SciQ datasets. PD, by measuring the performance difference between weak and strong models on samples, can more accurately reflect the relative difficulty and complexity of samples. In contrast, relying solely on metrics like PPL or educational value may not fully capture the differences in learning difficulty between samples.
@@ 3.3 Ablation Study
We focus on two main factors: (1) Size of the PD RM: We used PD calculated from various model combinations. (2) Choice of the RM: We use early and late checkpoints from a randomly trained 1.3B model to calculate PD and compare it with PD from RMs. The experimental results, as shown in Figure 4, demonstrate that regardless of the scale of the RM or the calculation method chosen, the results consistently outperform $Random$. This validates the robustness of our framework regarding PD calculation methods. Additionally, the PD calculated using the 100M-700M RMs slightly outperforms that of the early/end models, further supporting our hypothesis about the validity of approximating early and late checkpoints of the model with RMs, as they are comparable in terms of training FLOPS.
Figure 4: Ablation on PD calculation methods
@@
Table 3 presents the effects of the three preference functions in Section 2.3. Experimental results indicate that the S-Shape function outperforms other functions. Its slower initial decline compared to the linear function highlights the importance of starting with enough small PD data and gradually introducing large PD data to enhance performance. The Z-Shape function, which uses only small PD data initially and large PD data later, slightly outperforms the random setting. The linear, Z-shape and all S-shape parameter settings outperform $Random$, confirming the robustness of our framework.
@@
Z-shape - 55.6 23.6 86.2 33.8 68.9 53.6
Linear - 55.5 23.9 87.7 34.2 67.3 53.7
S-shape a=2.5 57.1 24.8 87.6 34.0 67.5 54.2
a=5.0 57.5 26.0 87.7 34.3 67.4 54.6
a=7.5 56.4 25.4 87.2 34.2 68.9 54.4
a=10.0 57.3 26.6 87.9 33.7 68.0 54.7
Table 3: Downstream tasks results for different preference functions. We report accuracy for each task, and the best performances are marked in bold. Abbreviations: H.P. = Hyper-parameters
@@ 3.4 Analysis
Training loss
We use 500M tokens from SlimPajama as our test set to compare test loss with the Random model on the 1.3B setting. Figure 5(a) shows that PDPC’s test loss initially declines slowly, then rapidly decreases, achieving a lower loss than Random by incorporating higher PD data later in training. The S-Shape function effectively minimizes loss, confirming optimal learning outcomes. Additionally, Figure 5(b) demonstrates that our method stabilizes the gradient norm, preventing oscillations and ensuring smoother model convergence. In the 3B model experiments (Figure 5(c) and 5(d)), the S-shape function significantly outperforms Random. It accelerates early loss reduction, speeds up convergence, and achieves a lower final loss, demonstrating superior performance. Additionally, it stabilizes the gradient norm.
@@ 3.5 Case study
Our analysis of 25M SlimPajama samples, sorted by PD values, shows significant distributional and semantic differences. High PD data mainly comes from Wikipedia and CommonCrawl, while low PD data is sourced from arXiv and GitHub, as seen in Figure 6. The significant performance improvement of our framework in later stages is largely attributed to the higher proportion of high-quality data from sources like Wikipedia, which primarily appears in the high PD data and is less prevalent in the low PD data.
@@
Semantic visualization using Sentence-T5 and t-SNE in Figure 7 illustrates data sampled in two ways: (a) uniformly from different PD partitions and (b) from extreme PD intervals after sorting. Uniform sampling results in high and low PD data being evenly distributed in semantic space, indicating semantic diversity. In contrast, extreme PD sampling leads to distinct semantic spaces, explaining the suboptimality of the PD sequential method due to induced data homogeneity, which can impede effective model training and generalization.
@@
We use raters across four QuRating dimensions to assess data quality under different PD values. Figure 8 shows consistent quality distributions in both high and low PD partitions, ensuring uniform quality throughout training and preventing the model from disproportionately learning from lower-quality data at any stage.
Figure 8: Data distribution across different quality dimensions.
@@
We evaluate the Spearman correlation coefficients between different PD types and find a strong correlation among PDs derived from RMs of varying sizes, which indicates that the ratio-based approach to computing PD maintains relative stability across samples. Calculating PD with smaller RMs produces results consistent with larger RMs, optimizing computational resources.
Furthermore, as illustrated in Figure 12 of Appendix C.2, larger size discrepancies among RMs result in broader PD distributions, which enhance data differentiation. This finding is supported by ablation tests, which showed that the 100M-1.3B PD calculations yielded the best results. Additionally, PD maintains a consistent distribution across domains. For instance, the PD between 100M and 700M models typically follows a normal distribution with a mean of 0.27. This stability makes PD a reliable metric, as it remains unaffected by changes in data properties. Consequently, organizing training data by PD ensures it does not skew towards specific sub-domains, allowing the model to encounter diverse data at every stage.
Semantic Feature Analysis
To explore differences between high and low PD data, we analyzed 1,000 randomly selected samples from each group using ten criteria focused on semantic features, including syntactic complexity, lexical diversity, and semantic depth. These criteria encompass polysemy, specialized terminology, cultural context, logical reasoning, humor, ethical dimensions, intricate sentence structures, scientific concepts, emotional nuances, and background knowledge. Each criterion is clearly defined for GPT-4o to assess with "yes" or "no" responses, as detailed in the Appendix C.4. Figure 10 indicates that PD is independent of other linguistic features, ensuring that reorganizing data by PD maintains diversity in these semantic properties.
Figure 10: Properties differences of high and low PD data.
@@ 4 Related Works
4.1 Curriculum Learning
Curriculum Learning (CL) is an adaptive training strategy that has gained attention in deep learning Forestier et al. (2022); Soviany et al. (2021); Ganesh and Corso (2020); Luo et al. (2020); Wang et al. (2022); Shi et al. (2024). It involves organizing training data to enable models to learn from simple to complex, enhancing efficiency and generalization.
CL strategies are widely applied in NLP Feng et al. (2023); Paulsen (2023); Li et al. (2024b). Collobert and Weston enhanced word embedding models using CL Collobert and Weston (2008). Recent advancements include Platanios et al.’s competence-based curriculum for neural machine translation, which reduces training time and improves translation quality Platanios et al. (2019), and Xu et al.’s application in natural language understanding, boosting performance and convergence Xu et al. (2020). In LLMs pretraining, CL is also applied Wu et al. (2024); Wettig et al. (2024).
4.2 Data Preprocessing for LLM Pretraining
In LLM pretraining, data preprocessing is crucial for ensuring dataset quality and relevance. Traditional methods apply expert-crafted rules to filter low-quality data Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024) and remove duplicates to maintain integrity Lee et al. (2022); Sorscher et al. (2022); Abbas et al. (2023); Soboleva et al. (2023); Tirumala et al. (2024). Although superficially effective, these techniques may not fully address the semantic richness needed for comprehensive training.
To enhance semantic content selection, some approaches use target data sources or proxy models for dataset curation Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023); Thakkar et al. (2023); Engstrom et al. (2024); Yu et al. (2024). Additionally, training classifiers for automated data selection is gaining traction. For example, Du et al. (2022) used a logistic regression binary classifier to evaluate data quality, while other studies have developed sophisticated scoring mechanisms Zhang et al. (2024b); Sachdeva et al. (2024). QuRating Wettig et al. (2024) illustrates a nuanced approach by training multiple raters to assess data contributions from various dimensions. Despite these advancements, current methods neglect the potential impact of data organization and sequencing on pretraining efficacy, highlighting an area for further exploration.
@@ 5 Conclusion
In this paper, we introduce the PDPC framework to optimize LLMs by dynamically adjusting pretraining data distribution. PD identifies challenging samples for weak models, scheduling them for later training stages. PDPC perceives and utilizes data preferred by LLMs, ensuring uninterrupted and continuous training while serving as the final data preprocessing step. Experiments demonstrate significant performance improvements over baselines, with the 3B model achieving over 4.1% gains in average accuracy, underscoring PDPC’s effectiveness.
@@ 6 Limitations and Future Works
We determined the S-shaped learning curve through functional exploration and used it as the basis for arranging the data sequence to train the model. In fact, we can also start from the newly trained model, re-explore new learning curves, and iteratively update our curriculum learning path. Optimizing the learning curve through multiple iterations could be one of our future research directions.
Appendix A Social Impact
@@ Appendix B Preliminary Exploration of Iterative Optimization of Preference Functions
When employing a grid search methodology, the size of the solution space scales as $n^{T}$, where $T$ represents the number of training stage divisions. Consequently, an increase in the number of domains $n$ results in an exponential expansion of the solution space.
In Section 2.3, we introduce a curriculum learning method that divides pretraining data into 2 domains and identifies the S-shape preference function through theoretical analysis and function search. This method is simple and efficient. However, in resource-rich scenarios, we also offer a more precise approach to simulate the model’s preferences at different pretraining stages. Specifically, after using the discovered preference function to guide the model’s pretraining, we conduct annealing experiments on checkpoints from different stages to explore the model’s preferences for data mixing ratios. Based on these preferences, we fit the model’s preference function to guide the next round of pretraining. This process is iterated until the model’s performance converges.

### v2

@@ 3.2 Main Results
Sequential PD High2Low 54.7$\downarrow$1.8 21.8$\downarrow$1.8 87.1$\uparrow$1.3 33.7$\downarrow$0.5 67.8$\uparrow$0.5 53.0$\downarrow$0.5
PD Low2High 56.1$\downarrow$0.4 21.3$\downarrow$2.3 86.2$\uparrow$0.4 34.4$\uparrow$0.2 67.6$\downarrow$0.2 53.1$\downarrow$0.4
PPL High2Low 45.5$\downarrow$11.0 20.6$\downarrow$3.0 71.2$\downarrow$14.6 30.3$\downarrow$3.9 63.7$\downarrow$3.6 46.3$\downarrow$7.2
PPL Low2High 47.8$\downarrow$8.7 17.9$\downarrow$5.7 72.7$\downarrow$13.1 29.1$\downarrow$5.1 62.4$\downarrow$4.9 46.0$\downarrow$7.5
Qu.Edu High2Low 57.2$\uparrow$0.7 26.4$\uparrow$2.8 85.4$\downarrow$0.4 33.0$\downarrow$1.2 66.2$\downarrow$1.1 53.6$\uparrow$0.1
Qu.Edu Low2High 56.8$\uparrow$0.3 26.0$\uparrow$2.4 84.1$\downarrow$1.7 33.5$\downarrow$0.7 67.9$\uparrow$0.6 53.7$\uparrow$0.2
Preference CL PPL S.R. 56.1$\downarrow$0.4 24.1$\uparrow$0.5 87.8$\uparrow$2.0 33.9$\downarrow$0.3 67.4$\uparrow$0.1 53.9$\uparrow$0.4
PPL S. 56.1$\downarrow$0.4 22.6$\downarrow$1.0 85.5$\downarrow$0.3 34.2   0.0 67.5$\uparrow$0.2 53.2$\downarrow$0.3
Qu.Edu S.R 56.7$\uparrow$0.2 24.9$\uparrow$1.3 86.2$\uparrow$0.4 33.6$\downarrow$0.6 66.9$\downarrow$0.4 53.7$\uparrow$0.2
Qu.Edu S. 55.5$\downarrow$1.0 24.8$\uparrow$1.2 87.8$\uparrow$2.0 34.0$\uparrow$0.2 67.4$\uparrow$0.1 53.9$\uparrow$0.4
PD S.R. 56.7$\uparrow$0.2 24.9$\uparrow$1.3 86.2$\uparrow$0.4 33.6$\downarrow$0.6 67.4$\uparrow$0.1 53.8$\uparrow$0.3
PDPC PD S. 57.3$\uparrow$0.8 26.6$\uparrow$2.9 87.9$\uparrow$2.1 33.7$\downarrow$0.5 68.0$\uparrow$0.7 54.7$\uparrow$1.2
Table 1: Downstream tasks results on 1.3B models with 100B tokens. We report accuracy for each task, and the best performances are marked in bold. Abbreviations: HellaSw. = HellaSwag, AVG. = Average, S.=S-shape Function, S.R.=S-shape Reverse Function.
@@
PDPC PD S. 69.7$\uparrow$1.1 35.8$\uparrow$2.1 95.3$\uparrow$0.7 76.3$\uparrow$0.3 35.8$\uparrow$8.1 35.6$\uparrow$8.1 36.1$\uparrow$8.9 54.9$\uparrow$4.1
Table 2: Downstream tasks results for different settings after training 3B models on 1T tokens.
@@
PDPC with $n=2$ consistently outperforms all baselines, regardless of the metric used, showing significant performance and convergence improvements over the baseline. Notably, the 3B model trained on 1T tokens with PDPC demonstrates an average accuracy increase of 4.1% across all benchmarks and 8.1% across MMLU and CMMLU, highlighting the effectiveness of our framework. Figure 3 depicts performance improvements in the 1.3B and 3B models as training progresses, with our method significantly outperforming Random in the latter half of pretraining. In this phase, data dominated by high PD is crucial, especially in the 3B model, highlighting the effectiveness of transitioning from low to high PD data, which significantly boosts model performance and promotes emergent capabilities.
@@
Sorting pretraining data by PD from low to high Sequential-PD-Low2High can outperform Random but falls short of the PDPC-PD-S. strategy. This limitation arises because sorting strictly by PD reduces data diversity, leading to homogeneity that restricts the model’s ability to handle complex tasks. In extreme cases, if the dataset contains duplicate samples, a complete sort would likely place identical samples in the same batch, which is detrimental to improving pretraining efficiency.
@@
Comparison of the results from Preference CL-PPL-S., Preference CL-Qu.Edu-S., and PDPC-PD-S. shows that using PD as a metric yields the best results, particularly on the ARC-C and SciQ datasets. PD can accurately reflect the relative difficulty and complexity of samples, which aligns well with the CL principle discussed in Section 1. In contrast, relying solely on PPL or educational value may not effectively capture the differences in sample difficulty required for this CL principle.
@@ 3.3 Ablation Study
We focus on two main factors: (1) Size of the RM: We use PD calculated from various model combinations. (2) Choice of the RM: We use early and late checkpoints from a randomly trained 1.3B model to calculate PD and compare it with PD from RMs. The experimental results, as shown in Figure 4, demonstrate that regardless of the scale of the RM or the calculation method chosen, the results consistently outperform Random. This validates the robustness of our framework regarding PD calculation methods. Additionally, PD from the 100M-700M RMs slightly outperforms that from the early/end models, further supporting our hypothesis that approximating the early and late checkpoints of the model with RMs is valid, as they are comparable in terms of training FLOPS.
Figure 4: Ablation on PD calculation methods.
@@
Table 3 presents the effects of the three preference functions in Section 2.3. Experimental results indicate that the S-shape function outperforms other functions. Its slower initial decline compared to the linear function highlights the importance of starting with enough low-PD data and gradually introducing high-PD data to enhance performance. The Z-shape function, which uses only low-PD data initially and high-PD data later, slightly outperforms Random. The linear, Z-shape, and all S-shape parameter settings outperform Random, confirming the robustness of our framework.
@@
Z-shape - 55.6$\downarrow$0.9 23.6   0.0 86.2$\uparrow$0.4 33.8$\downarrow$0.4 68.9$\uparrow$1.6 53.6$\uparrow$0.1
Linear - 55.5$\downarrow$1.0 23.9$\uparrow$0.3 87.7$\uparrow$1.9 34.2   0.0 67.3   0.0 53.7$\uparrow$0.2
S-shape a=2.5 57.1$\uparrow$0.6 24.8$\uparrow$1.2 87.6$\uparrow$1.8 34.0$\downarrow$0.2 67.5$\uparrow$0.2 54.2$\uparrow$0.7
a=5.0 57.5$\uparrow$1.0 26.0$\uparrow$2.4 87.7$\uparrow$1.9 34.3$\uparrow$0.1 67.4$\downarrow$0.1 54.6$\uparrow$1.1
a=7.5 56.4$\downarrow$0.1 25.4$\uparrow$1.8 87.2$\uparrow$1.4 34.2   0.0 68.9$\uparrow$1.6 54.4$\uparrow$0.9
a=10.0 57.3$\uparrow$0.8 26.6$\uparrow$3.0 87.9$\uparrow$2.1 33.7$\downarrow$0.5 68.0$\uparrow$0.7 54.7$\uparrow$1.2
Table 3: Downstream tasks results for different preference functions. We report accuracy for each task, and the best performances are marked in bold. Abbreviations: H.P. = Hyper-parameters.
@@ 3.4 Analysis
Loss Analysis
We sample 500M tokens from SlimPajama as test set to compare test loss with $Random$ on the 1.3B setting. Figure 5(a) shows that PDPC’s test loss initially declines slowly, then rapidly decreases, achieving a lower loss than $Random$ by incorporating higher PD data later in training. The S-shape function effectively helps to minimize loss. Additionally, Figure 5(b) demonstrates that our method stabilizes the gradient norm, ensuring smoother model convergence. In the 3B model experiments (Figure 5(c) and 5(d)), the S-shape function significantly outperforms $Random$. It accelerates early loss reduction, speeds up convergence, and achieves a lower final loss.
@@ 3.5 Case study
Our analysis of 25M texts from Slimpajama Soboleva et al. (2023) reveals significant differences in distribution and semantics between high-PD and low-PD data. High-PD data mainly comes from Wikipedia and CommonCrawl, while low-PD data is sourced from arXiv and GitHub, as seen in Figure 6. The significant performance improvement of PDPC in later stages is largely attributed to the higher proportion of high-quality data from sources like Wikipedia, which primarily appears in high-PD data and is less prevalent in low-PD data.
@@
To understand the semantic structure of the data, we use T5 Raffel et al. (2023) to generate dense vector representations of texts collected in two ways: (a) uniformly from different PD partitions, and (b) from extreme PD intervals (top/bottom 10%) after sorting. We then apply t-SNE for dimensionality reduction. Figure 7 illustrates the semantic visualization of the data points. Uniform sampling results in high and low PD data being evenly distributed in semantic space, indicating semantic diversity. In contrast, extreme PD sampling leads to distinct semantic spaces, explaining the suboptimality of Sequential-PD-Low2High, as it may result in a lack of data diversity in some batches during training, thereby affecting model performance.
@@
We use 4 raters from QuRatingWettig et al. (2024) to assess data quality. Figure 8 shows consistent quality distributions in both low-PD and high-PD parts, ensuring uniform quality throughout the pretraining process and preventing the model from learning from lower-quality data at any stage.
Figure 8: Distribution of low-PD and high-PD data across different quality dimensions.
@@
We evaluate the Spearman correlation coefficients between different PD types (detailed in Figure 11 of Appendix C.2) and find a strong correlation among PDs derived from RMs of varying sizes, which indicates that PD is a relatively stable metric. Calculating PD with smaller RMs yields results consistent with larger RMs, saving computational resources. Furthermore, larger size discrepancies among RMs result in broader PD distributions, which enhance data differentiation (detailed in Figure 12 of Appendix C.3). This finding is supported by ablation tests, which show that PD calculations using models ranging from 100M to 1.3B yield the best results. Additionally, PD maintains a consistent distribution across domains. For instance, the PD between 100M and 700M models generally appears to follow a normal distribution with an approximate mean of 0.27. Partitioning and organizing data using PD ensures that the data at each training step does not skew towards specific sub-domains, allowing the model to encounter a diverse range of data throughout the entire training process.
Semantic Properties Analysis
To explore the semantic features of high-PD and low-PD data, we analyze 1,000 samples randomly sampled from each part using 10 criteria focused on semantic features. These criteria encompass polysemy, specialized terminology, cultural context, logical reasoning, humor, ethical dimensions, intricate sentence structures, scientific concepts, emotional nuances, and background knowledge. Each criterion is clearly defined for GPT-4o to assess with "yes" or "no" responses. More details about the prompt can be found in the Appendix C.5. Figure 9 shows that PD is independent of other linguistic features. Partitioning and organizing data with PD maintains diversity in semantic properties, ensuring that model performance is not restricted by data homogeneity.
Figure 9: Semantic properties differences of low-PD data and high-PD data.
@@ 4 Related Works
Data preprocessing is crucial in LLM pretraining, ensuring dataset quality and integrity. Traditional methods use expert-crafted rules to filter low-quality data and remove duplicates Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024). Enhanced approaches leverage target data sources or proxy models for curation Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023). Automated data selection using classifiers is gaining traction; for example, Du et al. (2022) employed logistic regression to evaluate data quality, and other studies have developed sophisticated scoring mechanisms Zhang et al. (2024b); Sachdeva et al. (2024). QuRating Wettig et al. (2024) uses multiple raters to assess data contributions. Curriculum Learning (CL) complements these efforts by organizing training data from simple to complex, improving learning efficiency and generalization Forestier et al. (2022); Soviany et al. (2021). In NLP, CL enhances models, such as in word embeddings Collobert and Weston (2008) and neural machine translation Platanios et al. (2019). Recently, CL’s application in LLM pretraining is also growing Wu et al. (2024).
@@ 5 Conclusion
In this paper, we propose PDPC to address the limitations of consistent data distribution in pretraining LLMs. PDPC perceives the models’ preferences and utilizes different, model-preferred data as the models’ capabilities improve to enhance their performance. We introduce PD as a data metric and incorporate the preference function based on PD to predict data preferences, enabling the offline organization of data and ensuring uninterrupted pretraining. Experiments show that PDPC significantly outperforms the baselines, with the 3B model achieving an average improvement of 8.1% over Random on MMLU and CMMLU.
@@ 6 Limitations and Future Works
We determine the S-shaped learning curve through functional exploration and use it as the basis for arranging the data sequence to train the model. In fact, we can also start from the newly trained model, re-explore new learning curves, and iteratively update our curriculum learning path. Optimizing the learning curve through multiple iterations could be one of our future research directions.
Appendix A Ethical Considerations
@@ Appendix B Preliminary Exploration of Iterative Optimization of Preference Functions
When employing a grid search methodology, the size of the solution space scales as $n^{T}$, where $T$ represents the number of training steps. Consequently, an increase in the number of parts $n$ results in an exponential expansion of the solution space.
In Section 2.3, we introduce a curriculum learning method that partitions pretraining data into 2 parts and identifies the S-shape preference function through theoretical analysis. This method is simple and efficient. However, in resource-rich scenarios, we also offer a more precise approach to simulate the model’s preferences at different pretraining stages, as shown in Algorithm 2. Specifically, after using the discovered preference function to guide the model’s pretraining, we conduct annealing experiments on checkpoints from different stages to explore the model’s preferences for data mixing ratios. Based on these preferences, we fit the model’s preference function to guide the next round of pretraining. This process is iterated until the model’s performance converges.